OpenStudy (anonymous):

What mistake did I make?? Line 1: -2(x + 1) - 3 >= 11 Line 2: -2x + -2 - 3 >= 11 Line 3: -2x + -5 >= 11 Line 4: -2x <= 16 Line 5: x <= -8

OpenStudy (anonymous):

Well I do not know why it says "more than equals" >=

OpenStudy (anonymous):

So thats what I did wrong?

OpenStudy (anonymous):

I guess

OpenStudy (anonymous):

My teacher is asking me to write an essay on what I did wrong but I have no idea what I did wrong, so I have to be sure I know when I think I do

OpenStudy (jdoe0001):

do you know what to do to the inequality sign if you multiply by a negative value?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I think you're always supposed to flip the signs. Thats what I was taught. The little bit that my teacher taught us

OpenStudy (anonymous):

- Wasnt much

OpenStudy (anonymous):

oh that is what you mean nvm

OpenStudy (jdoe0001):

right.. so \(\bf Line 1:\ -2(x + 1) - 3 \ge 11\\ Line 2:\ -2x + -2 - 3 \ge 11\\ Line 3:\ -2x + -5 \ge 11\\ Line 4:\ {\large -2x {\color{red}{ \le}} 16}\quad \textit{did you multiply by a negative value?}\\ Line 5:\ x {\color{red}{ \le}} -8\)

OpenStudy (anonymous):

OH is that what sign you were trying to make IhateMathDotCom

OpenStudy (anonymous):

Yeah, it is. I looked up the sign, and I guess thats how you write it in computer text form. Like 10^5 to get 10 to the fifth power

OpenStudy (anonymous):

Oh sorry than listen to jdoe0001